You run at 6 mph. Your Dad leaves 15 min after you and runs at 9 mph. How long before he catches you? How far have each of you run?
Answer by mananth(16949)

You run at 6 mph. Your Dad leaves 15 min after you and runs at 9 mph. How long before he catches you? How far have each of you run?
..
you 6mph
dad 9mph
...
in 1/4 hour you run 1/4 *6 =3/2 miles
...
since both are running in same direction
the effective speed = 9-6 = 3mph.
time = d/r
t=3/2 / 3
t=3/2 *1/3
t= 1/2 hour
dad catches up after 1/2 hour
...
in 1/2 hour you have run 1/2 * 6 = 3 miles
dad has run 1/2 * 9 = 4.5 miles
m.ananth@hotmail.ca
...